a total of 4000 is invested:part at 11% the other at 13% . How much is invested at each if there is an annual rate of 450
Answer by richwmiller(17219)

We know the total amount of money invested. $4000
x+y=4000,
We know that the total interest for the year for the two accounts is $450
0.11*x+0.13*y=450
x=4000-y
We substitute for x
0.11*(4000-y)+0.13*y=450
We multiply out
440-0.11y+0.13*y=450
We combine like terms.
0.02*y=10
Isolate y
y=500 at 13%
x=4000-y
Calculate x
x=3500 at 11%
We check
0.11*3500+0.13*500=450
385+65=450
450=450
Since this statement is TRUE and neither amount is negative then it is ok